OpenStudy (anonymous):

Find the limit: lim x->2+ e^(3/(2-x))

OpenStudy (solomonzelman):

is it, \(\color{#000000 }{ \displaystyle \lim_{x\to2^+}e^{3/(2-x)} }\) ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

I will ask you the following question. \(\color{#000000 }{ \displaystyle \lim_{x\to2^+} \frac{1}{2-x}=? }\)

OpenStudy (anonymous):

- infinity?

OpenStudy (solomonzelman):

Yes, that is correct!

OpenStudy (solomonzelman):

\(\color{#000000 }{\LARGE \displaystyle \lim_{x\to2^+}e^{3/(2-x)} =e^{\lim_{x\to2^+}~3/(2-x)} }\)

OpenStudy (solomonzelman):

Well, technically, not, but you can think of that this way

OpenStudy (solomonzelman):

e^(-∞)=1/e^∞=0

OpenStudy (solomonzelman):

But, technically not so right, because we don't raise numbers to infinities.

OpenStudy (anonymous):

Okay so my answer would be zero right?

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

\(\color{#000000 }{\large \displaystyle y=\lim_{x\to2^+}e^{3/(2-x)} }\) \(\color{#000000 }{\large \displaystyle \ln y= \ln\left(\lim_{x\to2^+}e^{3/(2-x)}\right) }\) \(\color{#000000 }{\large \displaystyle \ln y= \left(\lim_{x\to2^+}\ln e^{3/(2-x)}\right) }\) \(\color{#000000 }{\large \displaystyle \ln y= \left(\lim_{x\to2^+}{3/(2-x)}\right) }\) So, if ln(y) tends to negative infinity, then y is 0.

OpenStudy (anonymous):

oh okay I see that makes more sense, I knew a ln was needed prove the answer, I just wasn't sure how to

OpenStudy (solomonzelman):

Well, you can just tell that 3/(2-x) will tend to -∞, and thus e^(-∞) = blah blah blah as I showed